OpenStudy (anonymous):

Solve for x: 3|x − 3| + 2 = 14 No solutions x = −1, x = 8.3 x = 0, x = 7 x = −1, x = 7

OpenStudy (anonymous):

I'm starting to do that in math YAY!

OpenStudy (anonymous):

First you have to get the absolute value by itself on one side. 3|x-3| + 2 = 14 3|x-3| = 12 |x-3| = 4 Then you solve for the negative and positive of the right side. x-3=4 x=7 x-3=-4 x=-1 So x = 7 or -1